# A Geometric View of Data Complexity: Efficient Local Intrinsic Dimension Estimation with Diffusion Models

### v1

@@ 3.3 FLIPD: A Fokker-Planck-Based LID Estimator
The LIDL estimator with DMs presented in subsection 3.2 provides a massive speedup over the naïve approach of training $m$ DMs, and over the method from subsection 3.1 requiring $m$ ODE solves. Yet, solving Equation 14 involves computing the trace of the Jacobian of $\hat{s}$ multiple times within an ODE solver, which, as mentioned in subsection 2.1, remains expensive. In this section we present our LID estimator, FLIPD, which circumvents the need for an ODE solver altogether. Recall that LIDL is based on Equation 8, which justifies the regression. Differentiating this equation yields that $\partial\log\varrho(x,\delta_{0})/\partial\delta\approx\text{LID}(x)-D$ for negative enough $\delta_{0}$, meaning that Equation 13 directly provides the rate of change that the regression in LIDL aims to estimate, from which we get
@@
$\text{LID}(x)\approx D+\dfrac{\partial}{\partial\delta}\log\varrho(x,\delta_{0})\approx D+\nu\big{(}t(\delta_{0});\hat{s},x\big{)}\eqqcolon\text{FLIPD}(x,t_{0}),$ (15)
@@
where $t_{0}\coloneqq t(\delta_{0})$. In addition to not requiring an ODE solver, computing FLIPD requires setting only a single hyperparameter, whereas regression-based estimators require $m$. Since $\nu(t(\delta);\hat{s},x)$ depends on $\delta$ only through $t(\delta)$, we can directly set $t_{0}$ as the hyperparameter rather than $\delta_{0}$, which avoids the potentially cumbersome computation of $t(\delta_{0})=\lambda^{-1}(e^{\delta_{0}})$: instead of setting a suitably negative $\delta_{0}$, we set $t_{0}>0$ sufficiently close to 00. In Appendix A we include explicit formulas for $\text{FLIPD}(x,t_{0})$ for common DMs.
@@
Theorem 3.1.
@@
We hypothesize that this result can be extended to non-linear submanifolds since, intuitively, every manifold can be locally linearly approximated (by its tangent space) and $\varrho(x,\delta)$ is “local as $\delta\rightarrow-\infty$” in the sense that its dependence on the values $p(y,0)$ becomes negligible as $\delta\rightarrow-\infty$ (because $\mathcal{N}(0,e^{2\delta}I_{D})$ approaches a point mass at 00) when $y$ is not in a close enough neighbourhood of $x$. However, we leave generalizing our result to future work.
@@ 4 Experiments
Throughout our experiments, we use variance-preserving DMs, the most popular variant of DMs, and compatible with DDPMs. We also include a “dictionary” to translate between the score-based formulation under which we developed FLIPD, and DDPMs, in Appendix C. Our code is available at https://github.com/layer6ai-labs/flipd; see Appendix D.1 for our hyperparameter setup.
@@
FLIPD requires setting $t_{0}$ close to 00 (since all the theory holds in the $\delta\rightarrow-\infty$ regime). It is important to note that DMs fitted to low-dimensional manifolds are known to exhibit numerically unstable scores $s(\cdot,t_{0})$ as $t_{0}\searrow 0$ [62, 40, 39]. Our first set of experiments examines the effect of $t_{0}$ on $\text{FLIPD}(x,t_{0})$ by varying $t_{0}$ within the range $(0,1)$.
@@
Here, we summarize our synthetic experiments in Table 1 using two metrics of performance: the mean absolute error (MAE) between the predicted and true LID for individual datapoints; and the concordance index, which measures similarity in the rankings between the true LIDs and the estimated ones (note that this metric only makes sense when the dataset has variability in its ground truth LIDs, so we only report it for the appropriate entries in Table 1). We compare against the NB and LIDL estimators described in subsection 2.2, as well as two of the most performant model-free baselines: LPCA [21, 11] and ESS [29]. For the NB baseline, we use the exact same DM backbone as for FLIPD (since NB was designed for variance-exploding DMs, we use the adaptation to variance-preserving DMs used in [31], which produces extremely similar results), and for LIDL we use $8$ neural spline flows. In terms of MAE, we find that FLIPD tends to be the best model-based estimator, particularly as dimension increases. Although model-free baselines perform well in simplistic scenarios, they produce unreliable results as LID increases or more non-linearity is introduced in the data manifold. In terms of concordance index, FLIPD achieves perfect scores in all scenarios, meaning that even when its estimates are off, it always provides correct LID rankings. We include additional results in the appendices: in Appendix D.5 we ablate FLIPD, finding that using kneedle indeed helps, and that FLIPD also outperforms the efficient implementation of LIDL with DMs described in subsection 3.2 that uses an ODE solver. We notice that NB with the setting proposed in [57] consistently produces estimates that are almost equal to the ambient dimension; thus, in Appendix D.6 we also show how NB can be significantly improved upon by using kneedle, although it is still outperformed by FLIPD in many scenarios. In addition, in Tables 6 and 7 we compare against other model-free baselines such as MLE [37, 42] and FIS [1]. We also consider datasets with a single (uni-dimensional) manifold where the average LID estimate can be used to approximate the global intrinsic dimensionality. Our results in Table 8 demonstrate that although model-free baselines indeed accurately estimate the global intrinsic dimension, they perform poorly when focusing on their pointwise LID estimates.
Image experiments
@@
We first focus on the simple image datasets MNIST [35] and FMNIST [66]. We flatten the images and use the same MLP architecture as in our synthetic experiments. Despite using an MLP, our DMs can generate reasonable samples (Appendix E.1) and the FLIPD curve for both MNIST and FMNIST is shown in 4(a). The knee points are identified at $t_{0}=0.1$, resulting in average LID estimates of approximately $130$ and $170$, respectively. Evaluating LID estimates for image data is challenging due to the lack of ground truth. Although our LID estimates are higher than those in [47] and [10], our experiments (Table 8 of Appendix D.4) and findings in [60] and [57] show that model-free baselines underestimate LID of high-dimensional data, especially images.
@@
When moving to more complex image datasets, the simplistic MLP score network backbone fails to generate high-quality samples. Therefore, we replace it with state-of-the-art UNets [50, 65] (see Appendix E.2). As long as the DM accurately fits the data manifold, our theory should hold, regardless of the choice of backbone. Yet, we see that UNets do not produce a clear knee in the FLIPD curves (see curves in Figure 10 of Appendix E.1). We discuss why this might be the case in Appendix E.1, and from here on we simply set $t_{0}$ as a hyperparameter instead of using kneedle. Despite FLIPD curves lacking a knee when using UNets, we argue that our estimates remain valuable measures of complexity. We took random subsets of $4096$ images from each of FMNIST, MNIST, SVHN [45], and CIFAR10 [34], and sorted them according to their FLIPD estimates. We show the top and bottom 5 images for each dataset in 4(b), and include more samples in Appendix E.3. Our visualization shows that higher FLIPD estimates indeed correspond to images with more detail and texture, while lower estimates correspond to less complex ones. Additionally, we show in Appendix E.4 that using only $50$ Hutchinson samples to approximate the trace term in FLIPD is sufficient for small values of $t_{0}$: FLIPD not only visually corresponds to image complexity, but also required only $50$ Jacobian-vector-products, resulting in a significant speedup over all prior work.
@@
Table 2: Spearman’s correlation between LID estimates and PNG compression size.
@@
Further, we quantitatively assess our estimates by computing Spearman’s rank correlation coefficient between different LID estimators and PNG compression size, used as a proxy for complexity in the absence of ground truth. As shown in Table 2, FLIPD has a high correlation with PNG, whereas model-free estimators do not. We find that the NB estimator correlates slightly more with PNG on MNIST and CIFAR10, but significantly less in FMNIST and SVHN. We also highlight that NB requires $\sim 3000$ function evaluations to produce a single LID estimate for MNIST and FMNIST, and $\sim 12000$ for SVHN and CIFAR10 – a massive increase in computation as compared to the $50$ Jacobian-vector-products we used for FLIPD. Moreover, in Appendix E.5, we analyze how increasing $t_{0}$ affects the estimator by re-computing the correlation with the PNG size at different values of $t_{0}\in(0,1)$. We see that as $t_{0}$ increases, the correlation with PNG decreases. Despite this decrease, we observe an interesting phenomenon: while characteristics of the image orderings change, qualitatively, the smallest $\text{FLIPD}(\cdot,t_{0})$ estimates still represent less complex data compared to the highest $\text{FLIPD}(\cdot,t_{0})$, even for relatively large $t_{0}$. We hypothesize that for larger $t_{0}$, similar to the “string within a doughnut” experiment in Figure 3, the orderings correspond to coarse-grained and semantic notions of complexity rather than fine-grained ones such as textures, concepts that a metric such as PNG size cannot capture.
(a) FLIPD of MNIST and FMNIST.
(b) Ordering small images.
Figure 4: Overview of image LID: (a) shows the FLIPD curves that are used to estimate average LID for MNIST and FMNIST; (b) compares images with small and large FLIPD estimates from FMNIST, MNIST, SVHN, and CIFAR10; and (c) compares LAION images with small and large FLIPD estimates using Stable Diffusion (top, $t_{0}=0.5$) and PNG compression sizes (bottom).
@@
Finally, we consider high-resolution images from LAION-Aesthetics [53] and, for the first time, estimate LID for extremely high-dimensional images with $D=3\times 512\times 512=786,\!432$. We use Stable Diffusion [49], a latent DM pretrained on LAION-5B [53]. This includes an encoder and a decoder trained to preserve relevant characteristics of the data manifold in latent representations. Since the encoder and decoder are continuous and effectively invert each other, we argue that the Stable Diffusion encoder can, for practical purposes, be considered a topological embedding of the LAION-5B dataset into its latent space of dimension $4\times 64\times 64=16,\!384$. Therefore, the dimension of the LAION-5B submanifold in latent space should be unchanged. We thus estimate image LIDs by carrying out FLIPD in the latent space of Stable Diffusion. Here, we set the Hutchinson sample count to $1$, meaning we only require a single Jacobian-vector-product. When we order a random subset of $1600$ samples according to their FLIPD at $t_{0}=0.3$, the more complex images are clustered at the end, while the least complex are clustered at the beginning: see Figure 1 and 4(c) for the lowest- and highest-LID images from this ordering, and Figure 25 in Appendix E.6 to view the entire subset and other values of $t_{0}$. In comparison to orderings according to PNG compression size (4(c)), FLIPD estimates prioritize semantic complexity over low-level details like colouration.
@@ 5 Conclusions, Limitations, and Future Work
In this work we have shown that the Fokker-Planck equation can be utilized for efficient LID estimation with any pre-trained DM. We have provided strong theoretical foundations and extensive benchmarks showing that FLIPD estimates accurately reflect data complexity. Although FLIPD produces excellent LID estimates on synthetic benchmarks, the lack of knees in FLIPD curves on image data when using state-of-the-art architectures is surprising, and results in unstable LID estimates which strongly depend on $t_{0}$. We see this behaviour as a limitation, even if FLIPD provides a meaningful measure of complexity in these cases. Given that FLIPD is tractable, differentiable, and compatible with any DM, we hope that it will find uses in applications where LID estimates have already proven helpful, including OOD detection, AI-generated data analysis, and adversarial example detection.
@@ B.2 Derivation of Equation 13
$\frac{\partial}{\partial\delta}\log\mathcal{N}^{\delta}_{d}(x^{\prime})=\frac{\partial}{\partial\delta}\left\{-\frac{d}{2}\log(2\pi)-d\delta-\frac{e^{-2\delta}}{2}\left\lVert x^{\prime}\right\rVert_{2}^{2}\right\}=-d+e^{-2\delta}\left\lVert x^{\prime}\right\rVert_{2}^{2}.$ (55)

### v2

@@ 3.3 FLIPD: An Efficient Fokker-Planck-Based LID Estimator
The LIDL estimator with DMs presented in subsection 3.2 provides a massive speedup over the naïve approach of training $m$ DMs, and over the method from subsection 3.1 requiring $m$ ODE solves. Yet, solving Equation 14 involves computing the trace of the Jacobian of $\hat{s}$ multiple times within an ODE solver, which, as mentioned in subsection 2.1, remains expensive. In this section we present our LID estimator, FLIPD, which circumvents the need for an ODE solver altogether. Recall that LIDL is based on Equation 8, which justifies the regression. Differentiating this equation yields that $\partial/\partial\delta\,\log\varrho(x,\delta_{0})\approx\text{LID}(x)-D$ for negative enough $\delta_{0}$, meaning that Equation 13 directly provides the rate of change that the regression in LIDL aims to estimate, from which we get
@@
$\text{LID}(x)\approx D+\dfrac{\partial}{\partial\delta}\log\varrho(x,\delta_{0})=D+\nu\big{(}t(\delta_{0});s,x\big{)}\approx D+\nu\big{(}t(\delta_{0});\hat{s},x\big{)}\eqqcolon\text{FLIPD}(x,t_{0}),$ (15)
@@
where $t_{0}\coloneqq t(\delta_{0})$. Computing FLIPD is very cheap since the trace of the Jacobian of $\hat{s}$ has to be evaluated only once when calculating $\nu(t(\delta_{0});\hat{s},x)$. As mentioned in subsection 2.1, exact evaluation of this trace has a cost of $\Theta(DF)$, but can be reduced to $\Theta(kF)$ when using the stochastic Hutchinson trace estimator with $k\ll D$ samples. Notably, FLIPD provides a massive speedup over the NB estimator – $\Theta(kF)$ vs. $\Theta(KF+D^{2}K)$ – especially in high dimensions where $K>D\gg k$. In addition to not requiring an ODE solver, computing FLIPD requires setting only a single hyperparameter, $\delta_{0}$. Furthermore, since $\nu(t(\delta);\hat{s},x)$ depends on $\delta$ only through $t(\delta)$, we can directly set $t_{0}$ as the hyperparameter rather than $\delta_{0}$, which avoids the potentially cumbersome computation of $t(\delta_{0})=\lambda^{-1}(e^{\delta_{0}})$: instead of setting a suitably negative $\delta_{0}$, we set $t_{0}>0$ sufficiently close to 00. In Appendix A we include explicit formulas for $\text{FLIPD}(x,t_{0})$ for common DMs.
@@
Theorem 3.1 (FLIPD Soundness: Linear Case).
@@
We conjecture that our theorem can be extended to non-linear submanifolds since it is a local result and every manifold can be locally linearly approximated by its tangent plane. More specifically, $\varrho(x,\delta)$ becomes “increasingly local as $\delta\rightarrow-\infty$” in the sense that its dependence on the values $p(x_{0},0)$ becomes negligible as $\delta\rightarrow-\infty$ when $x_{0}$ is not in a close enough neighbourhood of $x$; this is because $\mathcal{N}(0,e^{2\delta}I_{D})$ concentrates most of its mass around 00 as $\delta\rightarrow-\infty$ (see Equation 6). However, we leave generalizing our result to future work.
@@ 4 Experiments
Throughout this section, we use variance-preserving DMs, the most popular variant of DMs. We provide a “dictionary” to translate between the score-based formulation of FLIPD and DDPMs in Appendix C. We hope this will enable practitioners who are less focused on the theoretical aspects to effortlessly apply FLIPD to their pre-trained DMs. Our code is available at https://github.com/layer6ai-labs/flipd; see Appendix D for more experimental details.
4.1 Experiments on Synthetic Data
@@
FLIPD requires setting $t_{0}$ close to 00 since all the theory holds in the $\delta\rightarrow-\infty$ regime. It is important to note that DMs fitted to low-dimensional manifolds are known to exhibit numerically unstable scores $s(\cdot,t_{0})$ as $t_{0}\searrow 0$ [64, 41, 40]. This fact does not invalidate FLIPD, but it suggests that it might be sensitive to the choice of $t_{0}$. Our first set of experiments examines the effect of $t_{0}$ on $\text{FLIPD}(x,t_{0})$ by varying $t_{0}$ within the range $(0,1)$.
@@ Results
Here, we summarize our synthetic experiments in Table 1 using two metrics of performance: the mean absolute error (MAE) between the predicted and true LID for individual datapoints; and the concordance index, which measures similarity in the rankings between the true LIDs and the estimated ones (note that this metric only makes sense when the dataset has variability in its ground truth LIDs, so we only report it for the appropriate entries in Table 1). We compare against the NB and LIDL estimators described in subsection 2.2, as well as two of the most performant model-free baselines: LPCA [22, 12] and ESS [30]. For the NB baseline, we use the exact same DM backbone as for FLIPD (since NB was designed for variance-exploding DMs, we use the adaptation to variance-preserving DMs used in [32], which produces extremely similar results), and for LIDL we use $8$ neural spline flows. In terms of MAE, we find that FLIPD tends to be the best model-based estimator, particularly as dimension increases. Although model-free baselines perform well in simplistic scenarios, they produce unreliable results as LID increases or more non-linearity is introduced in the data manifold. In terms of concordance index, FLIPD achieves perfect scores in all scenarios, meaning that even when its estimates are off, it always provides correct LID rankings. We include additional results in the appendices: in Appendix D.5 we ablate FLIPD, finding that using kneedle indeed helps, and that FLIPD also outperforms the efficient implementation of LIDL with DMs described in subsection 3.2 that uses an ODE solver. We notice that NB with the setting proposed in [58] consistently produces estimates that are almost equal to the ambient dimension; thus, in Appendix D.6 we also show how NB can be significantly improved upon by using kneedle, although it is still outperformed by FLIPD in many scenarios. In addition, in Table 7 and Table 8 we compare against other model-free baselines such as MLE [38, 43] and FIS [1]. We also consider datasets with a single (uni-dimensional) manifold where the average LID estimate can be used to approximate the global intrinsic dimension. Our results in Table 9 demonstrate that although model-free baselines indeed accurately estimate global intrinsic dimension, they perform poorly when focusing on (pointwise) LID estimates.
4.2 Experiments with Fully-Connected Architectures on Image Data
@@
We first focus on the simple image datasets MNIST [36] and FMNIST [68]. We flatten the images and use the same MLP architecture as in our synthetic experiments. Despite using an MLP, our DMs can generate reasonable samples (Appendix E.1) and the FLIPD curve for both MNIST and FMNIST is shown in 4(a). The knee points are identified at $t_{0}=0.1$, resulting in average LID estimates of approximately $130$ and $170$, respectively. Evaluating LID estimates for image data is challenging due to the lack of ground truth. Although our LID estimates are higher than those in [48] and [10], our experiments (Table 9 of Appendix D.4) and the findings in [61] and [58] show that model-free baselines underestimate LID of high-dimensional data, especially images.
4.3 Experiments with UNet Architectures on Image Data
(a) FLIPD of MNIST and FMNIST.
(b) Ordering small images.
Figure 4: Overview of image LID: (a) shows the FLIPD curves that are used to estimate average LID for MNIST and FMNIST when using MLP backbones; (b) compares images with small and large FLIPD estimates from FMNIST, MNIST, SVHN, and CIFAR10 when using UNet backbones; and (c) compares LAION images with small and large FLIPD estimates using Stable Diffusion (top, $t_{0}=0.3$) and PNG compression sizes (bottom).
@@ 4.3 Experiments with UNet Architectures on Image Data
When moving to more complex image datasets, the backbone fails to generate high-quality samples. Therefore, we replace it with state-of-the-art UNets [51, 67] (see Appendix E.2). Surprisingly, we find that using kneedle with UNets fails to produce sensible LID estimates with FLIPD (see curves in Figure 10 of Appendix E.1). We discuss why this might be the case in Appendix E.1, and from here on we simply set $t_{0}$ as a hyperparameter instead of using kneedle. Although avoiding kneedle when using UNets results in increased sensitivity with respect to $t_{0}$, we argue that FLIPD remains a valuable measure of complexity as it produces sensible image rankings and it is highly correlated with with PNG compression length. We took random subsets of $4096$ images from each of FMNIST, MNIST, SVHN [46], and CIFAR10 [35], and sorted them according to their FLIPD estimates (obtained using UNet backbones). We show the top and bottom 5 images for each dataset in 4(b), and include more samples in Appendix E.3. Our visualization shows that higher FLIPD estimates indeed correspond to images with more detail and texture, while lower estimates correspond to less complex ones. Additionally, we show in Appendix E.4 that using only $k=50$ Hutchinson samples to approximate the trace term in FLIPD is sufficient for small values of $t_{0}$.
Further, we quantitatively assess our estimates by computing Spearman’s rank correlation coefficient between different LID estimators and PNG compression size, used as a proxy for complexity in the absence of ground truth. We highlight that although we expect this coefficient to be high, a perfect LID estimator need not achieve a correlation of $1$. As shown in Table 2, FLIPD has a high correlation
@@
Table 2: Spearman’s correlation between LID estimates and PNG compression size. FLIPD and NB were computed using the same UNet backbone.
@@
with PNG, whereas model-free estimators do not. We find that the NB estimator correlates slightly more with PNG on MNIST and CIFAR10, but significantly less in FMNIST and SVHN. Moreover, in Appendix E.5, we analyze how increasing $t_{0}$ affects FLIPD by re-computing the correlation with the PNG size at different values of $t_{0}\in(0,1)$. We see that as $t_{0}$ increases, the correlation with PNG decreases. Despite this decrease, we observe an interesting phenomenon: while the image orderings change, qualitatively, the smallest $\text{FLIPD}(\cdot,t_{0})$ estimates still represent less complex data compared to the highest $\text{FLIPD}(\cdot,t_{0})$, even for relatively large $t_{0}$. We hypothesize that for larger $t_{0}$, similar to the “string within a doughnut” experiment in Figure 3, the orderings correspond to coarse-grained and semantic notions of complexity rather than fine-grained ones such as textures, concepts that a metric such as PNG compression size cannot capture.
@@
We also consider high-resolution images from LAION-Aesthetics [54] and, for the first time, estimate LID for extremely high-dimensional images with $D=3\times 512\times 512=786,\!432$. We use Stable Diffusion [50], a latent DM pretrained on LAION-5B [54]. This includes an encoder and a decoder trained to preserve relevant characteristics of the data manifold in latent representations. Since the encoder and decoder are continuous and effectively invert each other, we argue that the Stable Diffusion encoder can, for practical purposes, be considered a topological embedding of the LAION-5B dataset into its latent space of dimension $4\times 64\times 64=16,\!384$. Therefore, the dimension of the LAION-5B submanifold in latent space should be unchanged. We leave an empirical verification of this hypothesis to future work and thus estimate image LIDs by carrying out FLIPD in the latent space of Stable Diffusion. Here, we set the Hutchinson sample count to $k=1$, meaning we only require a single Jacobian-vector-product. When we order a random subset of $1600$ samples according to their FLIPD at $t_{0}=0.3$, the more complex images are clustered at the end, while the least complex are clustered at the beginning: see Figure 1 and 4(c) for the lowest- and highest-LID images from this ordering, and Figure 25 in Appendix E.6 to view the entire subset and other values of $t_{0}$. In comparison to orderings according to PNG compression size (4(c)), FLIPD prioritizes semantic complexity over low-level details like colouration.
Table 3: Time, in seconds, to estimate LID for a single image.
Method MNIST/FMNIST SVHN/CIFAR10 LAION
FLIPD $\mathbf{0.10}$ $\mathbf{0.13}$ $\mathbf{0.20}$
NB $1.6$ $10.8$ $>9\times 10^{3}$
Finally, we compare the runtimes for computing FLIPD and NB for all models using UNet backbones. We show results in Table 3. For $28\times 28$ greyscale (MNIST/FMNIST) and $3\times 32\times 32$ low-resolution RGB (SVHN/CIFAR10) images, we use $k=50$ Hutchinson samples: at these dimensions, FLIPD achieves 10$\sim$ 10$\times$ 10 × and 100$\sim$ 100$\times$ 100 × respective speedups over NB. For $4\times 64\times 64$ LAION-Aesthetics images on the latent space of Stable Diffusion, we use $k=1$ for FLIPD, which ensures it remains highly tractable. At this resolution NB is completely intractable: constructing $S(x)$ for a single image $x$ takes $2.5$ hours, and computing NB would then still require performing a SVD on this $16,384\times 65,536$ matrix.
@@ 5 Conclusions, Limitations, and Future Work
In this work we have shown that the Fokker-Planck equation can be utilized for efficient LID estimation with any pre-trained DM. We have provided strong theoretical foundations and extensive benchmarks showing that FLIPD estimates accurately reflect data complexity. Although FLIPD produces excellent LID estimates on synthetic benchmarks, its instability with respect to the choice of network architecture on images is surprising, and results in LID estimates which strongly depend on this choice. We see this behaviour as a limitation, even if FLIPD nonetheless still provides a meaningful measure of complexity regardless of architecture. Given that FLIPD is tractable, differentiable, and compatible with any DM, we hope that it will find uses in applications where LID estimates have already proven helpful, including OOD detection, AI-generated data analysis, and adversarial example detection.
@@ B.2 Derivation of Equation 13
$\frac{\partial}{\partial\delta}\log\mathcal{N}^{\delta}_{d}(x^{\prime})=\frac{\partial}{\partial\delta}\left(-\frac{d}{2}\log(2\pi)-d\delta-\frac{e^{-2\delta}}{2}\left\lVert x^{\prime}\right\rVert_{2}^{2}\right)=-d+e^{-2\delta}\left\lVert x^{\prime}\right\rVert_{2}^{2}.$ (54)